OpenStudy (anonymous):

The perimeter of a square is 48 m. What is the length of a side and a diagonal?

Parth (parthkohli):

One side of the square would be 48/4 = 12 as all sides are equal and there are 4 sides. You may solve for the diagonal by applying the pythagorean theorem.  \( \color{Black}{\Rightarrow 12^2 + 12^2 = c^2}\)